In Excel, What does the YEARFRAC function do? This week we used the Date(), Year(), Month() & Day() functions. How do they work?
Answer #1

The Yearfrac function calculates and returns a decimal number which represents the fraction of the year between two dates.

Date function returns the date format of a number.We enter =DATE(2019,2,7) and the result will be 7-Feb-2019 depending upon the formatting applied.

Year function returns oly the year portion of a date. So if the date is enetered in cell A1 as above, we will enter = YEAR(A1) to get 2019 as the result.

Month and day functions will work to return the month and days of a given date.

=MONTH(A1) will return Feb

=DAY (A1) will return 7 in our example